14. The residual dividend model

The residual dividend policy approach to dividend policy is based on the theory that a firm’s optimal dividend distribution policy is a function of the firm’s target capital structure, the investment opportunities available to the firm, and the availability and cost of external capital. The firm makes distributions based on the residual earnings.

Consider the case of Yellow Duck Distribution Company:

Yellow Duck Distribution Company has generated earnings of $240,000,000. Its target capital structure consists of 60% equity and 40% debt. It plans to spend $86,000,000 on capital projects over the next year and expects to finance this investment in the same proportion as its capital structure. The company makes distributions in the form of dividends.

What will Yellow Duck Distribution Company’s dividend payout ratio be if it follows a residual dividend policy?

74.58%

78.50%

82.43%

86.35%

If Yellow Duck Distribution Company reduces the amount of its the forecasted capital budget, how will this affect the firm’s annual dividend, assuming that all other factors are held constant?

The amount that Yellow Duck Distribution Company will pay out in dividends this year will increase.

The amount that Yellow Duck Distribution Company will pay out in dividends this year will decrease.

Most firms have earnings that vary considerably from year to year and do not grow at a reliably constant pace. Furthermore, their required investment may change often. Which of these statements is the most accurate?

A residual dividend policy can’t be of any help to most firms.

Most firms can still use the concepts behind a residual dividend policy to make long-run decisions about dividends.

40% Debt 60% Equity
Answer #1

The income remaining after meeting all potential future investments are distributed as dividend under the residual dividend model

Capital Spending = $86,000,000

Equity portion = 86,000,000*60% = $51,600,000

Earnings = $240,000,000

Hence, dividends = 240,000,000-51,600,000 = $188,400,000

Payout ratio = Dividend/Earnings

= 188,400,000/240,000,000

= 78.50%

The amount that Yellow Duck Distribution Company will pay out in dividends this year will increase.

Since the number of funds required will reduce

Most firms can still use the concepts behind a residual dividend policy to make long-run decisions about dividends.


answered by: Goldener